OpenStudy (anonymous):

How can I find vertices on a parabola if I have the vertex, focus, directrix, and symmetry line?

OpenStudy (anonymous):

vertex: (1,-5) focus: (0,-5) Directrix: x=2 Symetry: y=-5

OpenStudy (anonymous):

\[-4x+4=y ^{2}+10y+25\]

OpenStudy (anonymous):

I get standard form: -4(x-1)=(y+5)squared

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Im trying to graph it but i think I should find some vertices to help graph it?

hero (hero):

The focus is \((0,-5)\). The directrix is \((2,y)\) So insert those points into the equation \((x - x_1)^2 + (y - y_ 1)^2 = (x - x_2)^2 + (y - y_2)^2\) to get: \((x - 0)^2 + (y - (-5))^2 = (x - 2)^2 + (y - y)^2\) Which simplifies to: \(x^2 + (y + 5)^2 = (x - 2)^2\) And expands to: \(x^2 + (y + 5)^2 = x^2 - 4x + 4\) \(x^2\) cancels on both sides to get: \((y + 5)^2= -4x + 4\) Now solving for x you get: \(-\dfrac{(y + 5)^2}{4} + 1 = x\)

hero (hero):

Once you graph it, it should look something like this: https://www.desmos.com/calculator/46lmaqxorx

OpenStudy (anonymous):



hero (hero):

Basically, if you have the focus and the directrix, you can find the parabola algebraically.

hero (hero):

If you look closely, you'll notice that what you have is the same as what I have.

OpenStudy (anonymous):

ok i think i see it lol i was just wondering because im drawing the parabola on graphing paper and I just have the focus and vertex...how do i eyeball the two lines?

OpenStudy (anonymous):

can I find some points on the actual parabola line?

hero (hero):

Of course. To do that, all you have to do is just create a table of values. Begin with y = {0,1,2...} then find the corresponding x values.

hero (hero):

It would be wise to use your calculator for this activity.

OpenStudy (anonymous):

i actually dont have a graphing calculator...

hero (hero):

There are online calculators you can use.

OpenStudy (anonymous):

can I plug points into the equation itself?

hero (hero):

You can, but that would be a long arduous process since it seems that many of the y values will be in fraction form.

hero (hero):

You need a way to be able to find the x values quickly.

hero (hero):

To make things easier, you should begin with the vertex of the parabola.

OpenStudy (anonymous):

so i got points like (0,-3) and (-5,0)

hero (hero):

I hope those are correct. Also don't forget the vertex.

OpenStudy (anonymous):

thanks!

hero (hero):

Mostly to graph it, all you need is the general direction of the graph.

hero (hero):

If you have three points, (which it seems you do), then that should be enough.

hero (hero):

Your (-5,0) seems to be an approximation rather than an exact point on the graph.

OpenStudy (anonymous):

I got the points from a graphing calculator, it seems right i dont know

OpenStudy (anonymous):

what does it mean when the vertex is at origin? is that (0,0)?

hero (hero):

The vertex is already given to you. It's (1,-5) remember?

OpenStudy (anonymous):

oh sorry this is a different problem im looking at...

hero (hero):

Yes, that's what it means. If it says "vertex at origin" then that corresponds to the point (0,0)

OpenStudy (anonymous):

thanks!